In a survey, 30 people were asked how much they spent on their child's last birthday gift. The results were roughly bell-shaped with a mean of $37 and standard deviation of $13. Construct a confidence interval at a 98% confidence level. Give your answers to one decimal place.
Answer #1

Solution :

Given that,

Point estimate = sample mean = \bar x = 37

sample standard deviation = s = 13

sample size = n = 30

Degrees of freedom = df = n - 1 = 29

At 98% confidence level the t is ,

\alpha = 1 - 98% = 1 - 0.98 = 0.02

\alpha / 2 = 0.02 / 2 = 0.01

t\alpha /2,df = t0.01,29 = 2.462

Margin of error = E = t\alpha/2,df * (s /\sqrtn)

= 2.462 * (13 / \sqrt 30)

= 5.8

The 98% confidence interval estimate of the population mean is,

\bar x - E < \mu < \bar x + E

37 - 5.8 < \mu < 37 + 5.8

31.2 < \mu < 42.8

(31.2, 42.8)